# An Efficient Framework for Crediting Data Contributors of Diffusion Models

### v2

@@ Abstract
As diffusion models are deployed in real-world settings, and their performance is driven by training data, appraising the contribution of data contributors is crucial to creating incentives for sharing quality data and to implementing policies for data compensation. Depending on the use case, model performance corresponds to various global properties of the distribution learned by a diffusion model (e.g., overall aesthetic quality). Hence, here we address the problem of attributing global properties of diffusion models to data contributors. The Shapley value provides a principled approach to valuation by uniquely satisfying game-theoretic axioms of fairness. However, estimating Shapley values for diffusion models is computationally impractical because it requires retraining on many training data subsets corresponding to different contributors and rerunning inference. We introduce a method to efficiently retrain and rerun inference for Shapley value estimation, by leveraging model pruning and fine-tuning. We evaluate the utility of our method with three use cases: (i) image quality for a DDPM trained on a CIFAR dataset, (ii) demographic diversity for an LDM trained on CelebA-HQ, and (iii) aesthetic quality for a Stable Diffusion model LoRA-finetuned on Post-Impressionist artworks. Our results empirically demonstrate that our framework can identify important data contributors across models’ global properties, outperforming existing attribution methods for diffusion models.
@@ 4.2 Experiment Settings
Sparsified fine-tuning. Magnitude-based pruning (Han et al., 2015) is used to remove model weights according to their magnitudes, resulting in sparse diffusion models with reduced parameters: from 35.7M to 19.8M for CIFAR-20 at a pruning ratio of 0.3, from 274M to 70.9M for CelebA-HQ at a ratio of 0.5, and from 5.1M to 2.6M for ArtBench (Post-Impressionism) at a ratio of 0.5. To estimate Shapley values, the pruned models are fine-tuned on data subsets corresponding to different contributor combinations, with 1,000 steps, 500 steps, and 200 steps for CIFAR-20, CelebA-HQ, and ArtBench (Post-Impressionism), respectively.
@@ 4.3 Baseline Methods
We mainly categorize baseline attribution methods into three categories [^6] TracIn (Pruthi et al., 2020) is not considered since intermediate checkpoints may not be available in practice.: (i) similarity-based methods (e.g., pixel similarity between the generated and training images); (ii) leave-one-out (LOO) and its approximate variants such as the influence functions (Koh & Liang, 2017) and TRAK-based methods (Park et al., 2023); and (iii) application-specific metrics such as the aesthetic score of each training image for ArtBench (Post-Impressionism). Because retraining diffusion models is computationally expensive, gradients are computed using the original model $\theta^{*}$ for TRAK-based methods, as done in Zheng et al. (2023). Local attribution scores are averaged for the images generated by the original model and used for computing global model properties. To aggregate datum-level attributions, the sum of attributions corresponding to each contributor is taken for the influence functions and TRAK-based methods, since this is the principled aggregation (Koh et al., 2019). For similarity-based and application-specific methods, datum-level attribution scores are aggregated for contributor attribution by taking either the average or maximum. More details about baseline methods can be found in Appendix E.
@@ 4.5 Experiment Results
In contrast, our approach, sparsified-fine-tuning (sparsified-FT) Shapley, computes contributor attribution using the Shapley value, achieving the highest LDS results of 61.48%, 26.34% , and 61.44% for CIFAR-20, CelebA-HQ, and ArtBench (Post-Impressionism), respectively. While leave-one-out (LOO) achieves 30.66% LDS on CIFAR-20, its performance declines as the number of contributors increase, e.g., CelebA-HQ and ArtBench (Post-Impressionism). This shows that attribution based on the marginal contribution of Shapley subsets with respect to $\mathcal{F}$ provides the most accurate importance score. Despite achieving the best results for CelebA-HQ compared to the baseline methods, we observe that the LDS performance of sparsified-FT Shapley, 26.32%, is relatively low compared to those of CIFAR-20 and ArtBench. We also perform evaluation across varying datamodel subset sizes ($\alpha=0.25,0.5,0.75$), and our approach consistently outperforms others (Appendix F).
@@
We analyze the top contributors identified by sparsified-FT Shapley in CIFAR-20, including contributors to classes such as motorcycles, buses, and lawnmowers (Figure 4). High-quality labeling by these contributors ensures that the training images contain clear and meaningful objects, which should result in low entropy (i.e., high confidence) of a classifier (i.e., Inception v3) (Barratt & Sharma, 2018). We therefore compute the entropy of each image and find that the top 20% classes have a lower average entropy of 5.18, compared to 6.03 for the remaining classes (Figure 16). For CelebA-HQ, the most important celebrities ranked by sparsified-FT Shapley encompass a diverse range of demographics and tend to belong to non-majority demographic clusters (Figure 16). Excluding images corresponding to these celebrities can have a negative impact on the diversity score. For ArtBench, we observe that images from the top contributors are more vivid and exhibit more vibrant colors, as shown in Figure 4.
@@ 5 Discussion
In this work, we introduce the problem of attributing global model properties to data contributors in the context of diffusion models. To address this problem, we develop an efficient framework that estimates the Shapley values for data contributors by leveraging model pruning and fine-tuning. This approach speeds up retraining and improves inference runtime, while ensuring accurate attribution. Our framework can have wide-ranging implications, such as creating incentives for data contributors, rewarding data labelers in a way that satisfies game-theoretic fairness, assessing labeling quality, and improving model performance and fairness. Empirical results across multiple datasets and global model properties show that our framework outperforms existing attribution approaches based on the diffusion loss, such as TRAK (Park et al., 2023) and its variant D-TRAK (Zheng et al., 2023), in crediting data contributors. There are several promising directions for future research. First, beyond fine-tuning, recent unlearning methods designed for diffusion models (Gandikota et al., 2023; Heng & Soh, 2023) could be explored in combination with various pruning strategies (Ding et al., 2019; Fang et al., 2024; Liu et al., 2021). Second, it can be useful to allocate contributor-level importance to datum-level importance for the data provided by each contributor, providing attribution at two levels of granularity. The Owen value, which modifies the Shapley value, is well suited for this scenario (Owen, 1977). Beyond diffusion models, our proposed framework can also be extended to models that are expensive to retrain, such as large language models (LLMs), whenever appropriate pruning and unlearning approaches are available. Finally, our approach and existing attribution methods for diffusion models assume access to training data and model parameters, which are not necessarily available for deployed models. For text-to-image models, membership inference and in-context learning may be useful in addressing this challenge (Hu & Pang, 2023; Wang et al., 2023b).
@@ F.4 Alternative Unlearning Approaches for Removal
In this section, we present the performance of unlearning approaches: fine-tuning (FT), gradient ascent (GA)Graves et al. (2021), and influence unlearning (IU)Izzo et al. (2021). To evaluate how well these approaches approximate retraining, we randomly selected 100 subsets from the Shapley sampling distribution, and we computed the FID score between the samples generated by the model retrained on each subset vs. models undergone by the unlearning methods. The resulting average FID scores are 25.5, 185.9, and 101.3 for sparsified fine-tuning (FT), sparsified GA, and sparsified IU, respectively (Table 12). In contrast, the average FID scores without sparsification are 43.1, 145.2, and 67.7 for fine-tuning (FT), GA, and IU, respectively (Table 12). These results indicate that while both GA and IU can effectively unlearn the target set, they significantly degrade overall image quality, Figure 14 and Figure 15. In contrast, sparsified FT demonstrates superior performance, producing better image quality than both GA and IU.

### v3

@@ Abstract
As diffusion models are deployed in real-world settings, and their performance is driven by training data, appraising the contribution of data contributors is crucial to creating incentives for sharing quality data and to implementing policies for data compensation. Depending on the use case, model performance corresponds to various global properties of the distribution learned by a diffusion model (e.g., overall aesthetic quality). Hence, here we address the problem of attributing global properties of diffusion models to data contributors. The Shapley value provides a principled approach to valuation by uniquely satisfying game-theoretic axioms of fairness. However, estimating Shapley values for diffusion models is computationally impractical because it requires retraining on many training data subsets corresponding to different contributors and rerunning inference. We introduce a method to efficiently retrain and rerun inference for Shapley value estimation, by leveraging model pruning and fine-tuning. We evaluate the utility of our method with three use cases: (i) image quality for a DDPM trained on a CIFAR dataset, (ii) demographic diversity for an LDM trained on CelebA-HQ, and (iii) aesthetic quality for a Stable Diffusion model LoRA-finetuned on Post-Impressionist artworks. Our results empirically demonstrate that our framework can identify important data contributors across models’ global properties, outperforming existing attribution methods for diffusion models [^1] Our code is available at https://github.com/q8888620002/data_attribution.
@@ 4.2 Experiment Settings
Sparsified fine-tuning. Magnitude-based pruning (Han et al., 2015) is used to remove model weights according to their magnitudes, resulting in sparse diffusion models with reduced parameters: from 35.7M to 19.8M for CIFAR-20, from 274M to 70.9M for CelebA-HQ, and from 5.1M to 2.6M for ArtBench (Post-Impressionism). To estimate Shapley values, the pruned models are fine-tuned on data subsets corresponding to different contributor combinations, with 1,000 steps, 500 steps, and 200 steps for CIFAR-20, CelebA-HQ, and ArtBench (Post-Impressionism), respectively.
@@ 4.3 Baseline Methods
We mainly categorize baseline attribution methods into three categories [^7] TracIn (Pruthi et al., 2020) is not considered since intermediate checkpoints may not be available in practice.: (i) similarity-based methods (e.g., pixel similarity between the generated and training images); (ii) leave-one-out (LOO) and its approximate variants such as the influence functions (Koh & Liang, 2017) and TRAK-based methods (Park et al., 2023); and (iii) application-specific metrics such as the aesthetic score of each training image for ArtBench (Post-Impressionism). Because retraining diffusion models is computationally expensive, gradients are computed using the original model $\theta^{*}$ for TRAK-based methods, as done in Zheng et al. (2023). Local attribution scores are averaged across the images generated by the original model for computing global model properties. To aggregate datum-level attributions, the sum of attributions corresponding to each contributor is taken for the influence functions and TRAK-based methods, since this is the principled aggregation (Koh et al., 2019). For similarity-based and application-specific methods, datum-level attribution scores are aggregated for contributor attribution by taking either the average or maximum. More details about baseline methods can be found in Appendix E.
@@ 4.5 Experiment Results
In contrast, our approach, sparsified fine-tuning (sparsified-FT) Shapley, computes contributor attribution using the Shapley value, achieving the highest LDS results of 61.48%, 26.34% , and 61.44% for CIFAR-20, CelebA-HQ, and ArtBench (Post-Impressionism), respectively. While leave-one-out (LOO) achieves 30.66% LDS on CIFAR-20, its performance declines as the number of contributors increase, e.g., CelebA-HQ and ArtBench (Post-Impressionism). This shows that attribution based on the marginal contribution of Shapley subsets with respect to $\mathcal{F}$ provides the most accurate importance score. Despite achieving the best results for CelebA-HQ compared to the baseline methods, we observe that the LDS performance of sparsified-FT Shapley, 26.32%, is relatively low compared to those of CIFAR-20 and ArtBench. We also perform evaluation with additional datamodel subset sizes ($\alpha=0.25,0.75$), and our approach consistently outperforms others (Appendix F).
@@
We analyze the top contributors identified by sparsified-FT Shapley in CIFAR-20, including contributors to classes such as motorcycles, buses, and lawnmowers (Figure 4). High-quality labeling by these contributors should ensure that the training images contain clear and meaningful objects, which should result in low entropy (i.e., high confidence) of a classifier (i.e., Inception v3) (Barratt & Sharma, 2018). We therefore compute the entropy of each image and find that the top 20% classes have a lower average entropy of 5.18, compared to 6.03 for the remaining classes (Figure 16). For CelebA-HQ, the most important celebrities ranked by sparsified-FT Shapley encompass a diverse range of demographics and tend to belong to non-majority demographic clusters (Figure 16). Excluding images corresponding to these celebrities can have a negative impact on the diversity score. For ArtBench, we observe that images from the top contributors are more vivid and exhibit more vibrant colors, as shown in Figure 4.
@@ 5 Discussion
In this work, we introduce the problem of attributing global model properties to data contributors of diffusion models. We develop an efficient framework to estimate the Shapley values for data contributors by leveraging model pruning and fine-tuning, speeding up retraining and inference runtime while ensuring accurate attribution. Our framework can have a range of implications, such as creating incentives for data contributors, rewarding data labelers in a way that satisfies game-theoretic fairness, assessing label quality, and improving model performance and fairness. Empirical results for multiple datasets and global model properties show that our framework outperforms existing attribution approaches based on the diffusion loss, such as TRAK (Park et al., 2023) and its variant D-TRAK (Zheng et al., 2023), in crediting data contributors. There are several promising directions for future research. Beyond fine-tuning, recent unlearning methods designed for diffusion models (Gandikota et al., 2023; Heng & Soh, 2023) could be explored in combination with various pruning strategies (Ding et al., 2019; Fang et al., 2024; Liu et al., 2021). Beyond diffusion models, our proposed framework can also be extended to models that are expensive to retrain, such as large language models (LLMs), whenever appropriate pruning and unlearning approaches are available. Finally, our approach and existing attribution methods for diffusion models assume access to training data and model parameters, which are not necessarily available for deployed models. For text-to-image models, membership inference and in-context learning may be useful in addressing this challenge (Hu & Pang, 2023; Wang et al., 2023b).
@@ 6 Ethics Statement
7 Acknowledgements
We thank members of the Lee lab for providing feedback on this project. This work was funded by the National Science Foundation [DBI-1759487]; and the National Institutes of Health [R01 AG061132, R01 EB035934, and RF1 AG088824].
@@ F.4 Alternative Unlearning Approaches for Removal
In this section, we present the performance of unlearning approaches: fine-tuning (FT), gradient ascent (GA) (Graves et al., 2021), and influence unlearning (IU) (Izzo et al., 2021). To evaluate how well these approaches approximate retraining, we randomly selected 100 subsets from the Shapley sampling distribution, and we computed the FID score between the samples generated by the model retrained on each subset vs. models undergone by the unlearning methods. The resulting average FID scores are 25.5, 185.9, and 101.3 for sparsified fine-tuning (FT), sparsified GA, and sparsified IU, respectively (Table 12). In contrast, the average FID scores without sparsification are 43.1, 145.2, and 67.7 for fine-tuning (FT), GA, and IU, respectively (Table 12). These results indicate that while both GA and IU can effectively unlearn the target set, they significantly degrade overall image quality, Figure 14 and Figure 15. In contrast, sparsified FT demonstrates superior performance, producing better image quality than both GA and IU.